OpenStudy (ray_sihota):

8a-1/6a+8=3/4 having problems finding the variable.

OpenStudy (ray_sihota):

would my Lcd be 6a+8(4)

OpenStudy (anonymous):

\(\Large 8a-\frac{1}{6a}+8=\frac{3}{4}\\\Large a(8a-\frac{1}{6a}+8)=\frac{3}{4}a\\\Large 8a^2-\frac{1}{6}+8a-\frac{3a}{4}=0\\\Large 8a^2+\frac{29a}{4}-\frac{1}{6}=0\) Now, use Quadratic Formula we got: \(\huge \color{blue}{a=\frac{\frac{-29}{4}\pm\sqrt{(\frac{29}{4})^2+4\times8\times \frac{1}{6}}}{2\times8}}\)

OpenStudy (nikato):

or is your problem like this? \[\frac{ 8a-1 }{ 6a+8 }=\frac{ 3 }{ 4 }\]

OpenStudy (ray_sihota):

yes

OpenStudy (ray_sihota):

i got the lcd 6a+8(4) 4(8a-1)=3(6a=8) 32a-4=18a+24 14a=28 a=2

OpenStudy (nikato):

yup. thats is correct

OpenStudy (nikato):

although i dont know why u call it lcd. isnt it just cross multiply?

OpenStudy (ray_sihota):

that how we do it in class . Multiply the LCD with all the equations and cross out the common factor less chance of mistakes.

OpenStudy (nikato):

ok then... but anyways, your answer is correct

OpenStudy (ray_sihota):

thanks again

OpenStudy (nikato):

no problem